# SyncAnyone: Implicit Disentanglement via Progressive Self-Correction for Lip-Syncing in the wild

### v1

@@ 3.3 Stage 2: Mask-Free Correction
Through this end-to-end restoration process, the network autonomously learns to disentangle the facial region into a stable, preservable identity/background component and a dynamic, modifiable mouth component, all without any explicit spatial guidance. This design elegantly solves the mask dilemma, as the model’s learned receptive field for modification is data-driven and dynamic, directly learning what to change based on audio-visual consistency rather than a fixed boundary.
Despite the use of a background compositing module, our synthetic data inevitably contains minor artifacts, such as unnatural blending at the boundaries. However, our training strategy is designed to inherently address this. By using the synthetic data as the condition and the original, pristine data as the target, the model learns to automatically eliminate these imperfections. During inference, the input video frames are artifact-free, thus ensuring that the generation quality is not compromised.
Efficient Mask-Free Inference. The primary advantage of this implicit disentanglement paradigm is the exceptional simplicity and efficiency of its inference process. The Stage-1 model incurs significant computational overhead because concatenating multiple reference frames along the temporal dimension substantially increases the sequence length of the latents processed by the attention mechanism. In contrast, the Stage-2 model operates much more efficiently. It receives the full conditioning frame via the Y channel, eliminating the need for any additional temporal computations or reference frames at inference time. The model then directly and efficiently generates a video frame with the correct mouth shape and a lossless background, perfectly resolving the core issues of background consistency and inference efficiency.
4 Experiment
Table 1: Quantitative comparison with previous methods on HDTF Dataset.
HDTF Dataset
Method Full Reference Metrics No Reference Metrics Lip Sync
FID $\downarrow$ FVD $\downarrow$ CSIM $\uparrow$ NIQE $\downarrow$ BRISQUE $\downarrow$ HyperIQA $\uparrow$ LMD $\downarrow$ LSE-C $\uparrow$
Wav2Lip [prajwal2020lip] 14.912 543.340 0.852 6.495 53.372 45.822 10.007 7.630
VideoReTalking [cheng2022videoretalking] 11.868 379.518 0.786 6.333 50.722 48.476 8.848 7.180
TalkLip [wang2023seeing] 16.680 691.518 0.843 6.377 52.109 44.393 15.954 5.880
IP-LAP [zhong2023identity] 9.512 325.691 0.809 6.533 54.402 50.086 7.695 7.260
Diff2Lip [mukhopadhyay2024diff2lip] 12.079 461.341 0.869 6.261 49.361 48.869 18.986 7.140
MuseTalk [zhang2024musetalk] 8.759 231.418 0.862 5.824 46.003 55.397 8.701 6.890
LatentSync [li2024latentsync] 8.518 216.899 0.859 6.270 50.861 53.208 17.344 8.050
OminiSync [peng2025omnisyncuniversallipsynchronization] 7.855 199.627 0.875 5.481 37.917 56.356 7.097 7.309
Ours ($1^{st}$ stage) 8.245 223.594 0.877 5.738 39.748 54.912 7.089 7.776
Ours ($2^{nd}$ stage) 7.331 153.284 0.881 5.219 32.805 57.233 7.132 7.870
@@ 4.1 Experiment Settings
Datasets. We collected approximately 100 hours of multilingual speech data, which covers scenarios such as large pose angles, occlusions, and rapid background changes. We split these data by shots and used a face detector to extract the facial bounding boxes. Bilateral filtering was applied to these boxes to obtain smooth and compact tracking trajectories. During training, random jumps were introduced into the bounding boxes to simulate scene cuts during actual inference.
Implementation Details. We use the Wan2.1 [wan2025] 1.3B as our base model. All models are trained on 8 NVIDIA A100 GPUs with a total batch size of 8. The training process is divided into two distinct stages:
Stage 1: In the first stage, we train the mask-inpainting model $G_{inp}$ for 100k iterations. We use the AdamW optimizer [loshchilov2019decoupledweightdecayregularization] with $\beta_{1}=0.9,\beta_{2}=0.999$, and a weight decay of $0.01$. The learning rate is initialized to $2e^{-4}$ and follows a cosine annealing schedule with a warm-up of 5k iterations.
@@
4.2 Compared Methods
We benchmark our method against several state-of-the-art (SOTA) approaches that represent distinct technical paradigms. PersonaTalk [zhang2024personatalk], a GAN-based method, animates lip shapes by translating audio features into 3DMM BlendShape coefficients. LatentSync [li2024latentsync], a representative of Stable Diffusion-based techniques, inpaints the mouth region conditioned on audio after masking it in the driving video. Sharing the same DiT backbone as our method, OmniSync [peng2025omnisyncuniversallipsynchronization] achieves background-preserving lip-syncing by selectively applying a higher level of classifier free guidance to the oral region and a lower level to the rest of the frame. Given that the source codes of OmniSync and PersonaTalk are not open-sourced, we directly utilized their official provided services to generate the experimental results.
Figure 4: Lip-syncing results for different syllables.
4.3 Qualitative Comparison
Fig. 3 presents a comparative analysis of the processing results of our method against other schemes across diverse scenarios.
  • Normal Scenario: For frontal close-up scenes, our method achieves superior lip generation quality. Compared with GAN-based and SD-based schemes, our approach, built on the DiT backbone, facilitates the generation of high-quality results more effectively. Fig. 4 presents the driving results of our algorithm under different audio signals, demonstrating that our algorithm achieves high lip synchronization accuracy and lip clarity.
  • Extreme Pose: The second column illustrates the performance of different algorithms under large-pose angles. Our method produces side profiles with higher clarity, while the other methods exhibit obvious artifacts. This advantage stems from the fact that our algorithm does not rely on strict facial key point detection; instead, the model learns pose-specific processing strategies for side faces directly from the data. In contrast, GAN-based schemes fail to generate valid results due to their inability to recover the 3D facial structure from side views.
  • Occlusion: The third column demonstrates model performance under occluded conditions. It can be observed that our method preserves occluders completely while modifying lip movements. GAN-based methods generate lips with upward displacement and poor quality, as occlusions disrupt 3D facial detection. LatentSync, which employs mask inpainting, suffers from partial loss of occluders.
  • Scene Cut: The fourth column shows algorithm performance in scene cut scenarios. Benefiting from the data augmentation with manually synthesized cuts, our algorithm handles hard cuts robustly. OminiSync exhibits residual artifacts from the previous frame. PersonaTalk, being single-frame methods, are unaffected by cuts.
  • Stylish: Leveraging the inherent capabilities of the video backbone, our algorithm effectively drives lip movements in anime-style scenes. In contrast, PersonaTalk produce significant artifacts, and both LatentSync and OminiSync suffer from lip quality degradation.
  • Rapid Background Change: Owing to the mask-free training strategy, our algorithm modifies lip movements while preserving the original background even under high-frequency background variations. Due to the inability to access complete background information directly during inference, the background regions generated by OminiSync and LatentSync appear blurred.
4.4 Quantitative Results
To conduct a quantitative comparison with other methods, we evaluated the performance of our approach on the HDTF [zhang2021flow] dataset, and the results are presented in Tab. 1. Benefiting from the enhanced generation capability and background preservation ability enabled by the two-stage model, our algorithm achieves state-of-the-art (SOTA) results on distribution distance-related metrics such as FID and FVD. The leading performance on the CSIM metric also demonstrates that our algorithm possesses superior identity preservation capability. For image quality-related metrics, including NIQE, BRISQUE, and HyperIQA, our algorithm exhibits high performance consistently. Additionally, our method achieves performance comparable to SOTA approaches on lip motion-related metrics, namely LMD and LSE-C.
@@ 4.4 Quantitative Results
4.5 Ablation Study

### v3

@@ 4 Experiment
Figure 4: Lip-syncing results for different syllables.
@@ 4.1 Experiment Settings
Datasets.
We collected approximately 100 hours of multilingual speech data, which covers scenarios such as large pose angles, occlusions, and rapid background changes. We split these data by shots and used a face detector to extract the facial bounding boxes. Bilateral filtering was applied to these boxes to obtain smooth and compact tracking trajectories. During training, random jumps were introduced into the bounding boxes to simulate scene cuts during actual inference.
Implementation Details. We use the Wan2.1 (Wan et al., 2025) 1.3B as our base model. All models are trained on 8 NVIDIA A100 GPUs with a total batch size of 8. The training process is divided into two distinct stages:
Stage 1: In the first stage, we train the mask-inpainting model $G_{inp}$ for 100k iterations. We use the AdamW optimizer (Loshchilov and Hutter, 2019) with $\beta_{1}=0.9,\beta_{2}=0.999$, and a weight decay of $0.01$. The learning rate is initialized to $2e^{-4}$ and follows a cosine annealing schedule with a warm-up of 5k iterations.
@@
4.2 Experimental Comparisons
Our evaluation is structured into two parts: a rigorous quantitative benchmark on established datasets and a qualitative analysis focused on challenging, real-world scenarios.
Quantitative Evaluation. To ensure a fair and comprehensive comparison, we evaluate our method against a wide range of baselines representing different technical paradigms. We conduct these experiments on two popular open-source datasets: HDTF (Zhang et al., 2021) and VFHQ (Wang et al., 2022), which provide a diverse set of high-resolution talking head videos for benchmarking lip-sync accuracy and visual fidelity.
Qualitative Evaluation. To demonstrate the robustness of SyncAnyone in ”in-the-wild” scenarios, we perform qualitative comparisons on a curated set of challenging videos featuring extreme head poses, partial occlusions, rapid background dynamics, and stylish artistic content. In these evaluations, we primarily benchmark against the top-performing SOTA methods: OmniSync, PersonaTalk, and LatentSync, focusing on their ability to maintain identity and background consistency while performing precise lip editing.
Implementation Note. As the source codes for OmniSync and PersonaTalk are not publicly available, we utilized their official provided web services to generate all results for comparison, ensuring that we evaluate the best possible performance of these state-of-the-art models.
4.2.1 Quantitative Results
We evaluate SyncAnyone on randomly sampled subsets from the HDTF (Zhang et al., 2021) and VFHQ (Wang et al., 2022) datasets, following established protocols to ensure a fair comparison. As shown in Tab. 1, our Stage 2 model consistently achieves state-of-the-art results across visual quality, temporal stability, and identity preservation metrics. Notably, the substantial lead in FID and FVD scores highlights our model’s ability to generate high-fidelity frames with exceptional temporal coherence. Unlike mask-inpainting methods such as LatentSync (Li et al., 2024) and MuseTalk (Zhang et al., 2024b), which often introduce flickering or blurring artifacts near the mask boundaries, our mask-free paradigm ensures pixel-perfect background preservation and superior identity consistency, as reflected in the leading CSIM scores. In terms of synchronization, SyncAnyone maintains highly competitive LSE-C and LMD values, demonstrating that our implicit disentanglement mechanism successfully maintains precise audio-visual alignment without relying on explicit spatial constraints. Furthermore, the significant performance leap from Stage 1 to Stage 2 validates the effectiveness of our Progressive Self-Correction (PSC) framework. While Stage 1 provides robust motion priors, Stage 2 effectively eliminates mask-induced synthetic artifacts and refines high-frequency facial details. This transition results in a more realistic and stable final output, striking an ideal balance between generative realism and precise lip-synchronization.
@@ 4.2.1 Quantitative Results
4.2.2 Qualitative Comparison
Fig. 3 illustrates a comparative analysis between our method and state-of-the-art baselines across several challenging scenarios.
Normal Scenario: In frontal close-ups, our DiT-based approach produces superior lip clarity and synchronization compared to GAN and SD-based methods. As shown in Fig. 4, SyncAnyone accurately captures diverse phonetic articulations with high visual fidelity.
Extreme Pose: At large pose angles, our method maintains profile integrity and sharp details, whereas landmark-reliant GANs often fail to recover valid 3D structures. This advantage stems from our data-driven strategy, which learns pose-specific facial dynamics directly without explicit facial priors.
Occlusion: Our method effectively preserves occluding objects (e.g., hands or microphones) while modifying lip movements. In contrast, GAN-based methods often generate displaced mouths due to detection failures, and inpainting-based methods like LatentSync suffer from partial occluder erosion.
Scene Cut: Benefiting from cut-aware data augmentation, SyncAnyone handles hard scene cuts robustly. While OmniSync often exhibits residual ”ghosting” artifacts from preceding frames, our model ensures clean transitions and temporal stability.
Stylish: Leveraging the strong generative priors of the video transformer backbone, our model generalizes effectively to anime and artistic styles. Baselines, however, typically suffer from significant quality degradation and unnatural artifacts in these non-photorealistic domains.
Rapid Background Change: Thanks to our mask-free training strategy, our model achieves pixel-perfect background consistency even under high-frequency variations. Conversely, methods like OmniSync and LatentSync often produce blurred backgrounds as they fail to access or preserve complete background information during inference.
4.2.3 Ablation Study